# Structural Properties of Entropic Vectors and Stability of the Ingleton Inequality

### v1

@@ 1 Introduction
In this paper we study inequalities for Shannon entropy, and, in particular, variants of the Ingleton inequality in the entropic setting. It is well known that the Ingleton inequality in its original form does not hold for entropies of certain tuples of jointly distributed random variables; see, e.g. [Mat95, Csi96, ZY98, Mat99, Ham+00]. However, under certain restrictions (or with suitable additional terms) the inequality may still hold. We discuss several manifestations of this important phenomenon.
We begin with a brief review of the standard framework for information inequalities. For a tuple of random variables $(X_{i}\;{\bm{:}}\;i=0,\dots,n-1)$ its entropy profile is the function on the power set of the set $\left\{0,\dots,n-1\right\}$ indexing the tuple. The value of this function on $J\in 2^{[n]}$ is the entropy of the joint distribution $(X_{i}\;{\bm{:}}\;i\in J)$. Since the entropy of the empty collection is always zero, we drop the corresponding value from consideration. Thus, every entropy profile is effectively a point in $\mathbb{R}^{2^{n}-1}$. In particular, for a four-tuple of random variables, its entropy profile is a 15-dimensional vector.
A point in $\mathbb{R}^{2^{n}-1}$ is called entropic if it is the entropy profile of the probability distribution of some tuple $(X_{i}\;{\bm{:}}\;i=0,\dots,n-1)$. The set of all entropy profiles for all $n$-tuples of random variable is usually denoted $\Gamma_{n}^{\star}$. The points in the closure of the set of entropy profiles are called almost entropic points. This set is denoted by $\bar{\Gamma}_{n}^{\star}$. It has been shown in [ZY97] that this set is in fact a convex cone. The geometry of this cone is quite non-trivial. It is known, for example, that for all $n\geq 4$ the set $\bar{\Gamma}_{n}^{\star}$ is not polyhedral, [Mat07]). An exact description of this cone for $n\geq 4$ is still unknown. However, there are some known lower and upper bounds for this set. In what follows we briefly discuss the basic facts about $\bar{\Gamma}_{4}^{\star}$.
@@
The set of all $\varphi$ satisfying these inequalities is traditionally denoted $\Gamma_{4}$ (the cone of submodular functions on a four-point set). This set is clearly a superset (i.e., an upper bound) of $\bar{\Gamma}_{4}^{\star}$.
Some sufficient conditions for being an entropic point can be formulated with the help of the Ingleton inequalities. In what follows we use the notation: for a set function $\varphi$ define
@@
With this notation, the Ingleton inequality can be written in the following compact form:
@@
Permuting the set $\left\{0,1,2,3\right\}$ we obtain $\binom{4}{2}=6$ similar inequalities that differ in the order of the variables. We denote by $\Gamma^{\rm Ing}_{4}$ the cone of all non-negative, monotone, submodular functions $\varphi$ on $\left\{0,1,2,3\right\}$ that satisfy, in addition to the Shannon inequalities, the six versions of the Ingleton inequality on a four-point set. This set is a subset (lower bound) of the set of almost entropic points: it is known that every $\varphi$ in $\Gamma^{\rm Ing}_{4}$ is almost entropic, see [MS95] (independently, this was proved in [Ham+00]).
@@
The sets $\Gamma_{4}$ and $\Gamma_{4}^{\mathrm{Ing}}$ are both polyhedral cones, and the set of entropic points is sandwiched between the two: $\Gamma_{4}^{\mathrm{Ing}}\subsetneq\bar{\Gamma}_{4}^{\star}\subsetneq\Gamma_{4}.$ The closure of $\Gamma_{4}\setminus\Gamma_{4}^{\mathrm{Ing}}$ is the union of six closed convex cones with disjoint interiors. They are symmetric under permutations of the set $\left\{0,1,2,3\right\}$. From now on we consider just one of these cones in which the strict version of inequality (1) is violated, and we refer to it as the non-Ingleton cone.
@@
The non-Ingleton cone has a 14-dimensional simplex as its base. One of the 13-dimensional faces of this simplex is defined by the Ingleton equality and is a common face with the base of $\Gamma_{4}^{\mathrm{Ing}}$. The remaining vertex of the simplex is the so-called special point. It is neither an entropic nor an almost entropic point, as was shown in [ZY98]. The study of the cone of almost entropic points essentially reduces to the question of what neighborhood around the special point is free of almost entropic points, or, equivalently, what neighborhood of the face shared between the Ingleton and non-Ingleton regions is filled with almost entropic points. [^1] If we denote by $\alpha_{1},\dots,\alpha_{15}$ the convex coordinates in the non-Ingleton simplex (the base simplex of non-Ingleton cone), such that $\alpha_{15}=1$ is the defining equation of the special point, while $\alpha_{15}=0$ is the equation of the face common with the Ingleton-region, then the almost-entropic region is described by the inequality $\alpha_{15}\leq\Phi(\alpha_{1},\dots,\alpha_{14})$ where $\Phi$ is some concave function on the $13$-dimensional simplex. Ultimately, the complete characterization of $\bar{\Gamma}^{*}_{4}$ reduces to the description of the function $\Phi$. The main results of this paper contribute to a better understanding of the behavior of $\Phi$ near the boundary of its domain. A detailed discussion of the geometry of the cones $\bar{\Gamma}_{4}^{\star}$, $\Gamma_{4}^{\mathrm{Ing}}$, and $\Gamma_{4}$, and of the non-Ingleton cone can be found, e.g., in [Csi16, Csi25] and in [MC16, MP19]. In general, the complete description of $\bar{\Gamma}_{4}^{\star}$ remains far from being understood.
Constraint versions of the Ingleton inequality.
@@
In what follows we discuss the Ingleton inequality for a four-tuple of random variables. To avoid ambiguity and improve the readability of the notation we denote the four-tuple of jointly distributed random variables $(X,Y,A,B)$, and rewrite (1) with the usual notation for the conditional mutual information as
@@
We will consider several cases in which some of the quantities
I(A:X|Y),I(A:Y|X),I(B:X|Y),I(B:Y|X),I(X:Y|A),I(X:Y|B)I(A:X|Y),I(A:Y|X),I(B:X|Y),I(B:Y|X),I(X:Y|A),I(X:Y|B) (2)
are small compared with other information quantities. It is known that (ING) holds for joint distributions $(X,Y,A,B)$ satisfying certain independence constraints, that is, whenever some of the quantities in (2) vanish. More specifically, the following implications are known:
@@
as well as the symmetric statements obtained by exchanging $A$ with $B$. These conditional constraints were stated explicitly in [KR13], although they were implicitly known earlier, since they follow directly from the infinite families of inequalities established by Matúš in his seminal paper [Mat07].
We focus on the possible violation of the Ingleton inequality in the regime where some of the mutual information terms in (2) are small but still positive. We consider three cases:
@@
$\max\left\{I(A:X|Y),I(A:Y|X)\right\}\leq\varepsilon\cdot I(X:Y).$
@@
$\max\left\{I(X:Y|A),I(A:Y|X)\right\}\leq\varepsilon\cdot I(X:A).$
@@
Proofs of Theorems A and B are given in Section 4, Theorem C is proven in Section 6.
@@
$\quad{}+\frac{1}{k}I(X:Y|A)+\frac{k+1}{2}\big(I(A:X|Y)+I(A:Y|X)\big)\quad\text{for}\ k\in\mathbb{N}$
@@
$\quad+\frac{1}{k}I(A:X|Y)+\frac{k+1}{2}\big(I(X:Y|A)+I(A:Y|X)\big)\quad\text{for }k\in\mathbb{N}$
proven by Matúš in [Mat07] (it suffices to choose a suitable value of $k$). In fact, our proof of the lemma above runs parallel to the argument of Matúš — we apply the Copy Lemma and the Ahlswede–Körner lemma, while Matúš used in [Mat07] a similar (and formally equivalent) method of adhesive extensions.
@@
This article presents more direct proofs of Theorems ABC without deducing preliminary infinite families of non-Shannon–type inequalities for entropy. This approach offers, in our view, several noteworthy advantages. To begin with, it is conceptually more intuitive, as it is grounded in certain structural properties of probability distributions and in the intuition of materializable common information, which we elaborate on below. This perspective enables us to reveal and prove properties that are stronger than a mere conditional form of the Ingleton inequality; see Theorem F below. In addition, the structural property we identify appears particularly amenable to computer-assisted proof systems. We believe that this may enable computer-assisted discoveries of new conditional information inequalities, in much the same spirit as non-Shannon–type inequalities and their applications have been established in [DFZ11] and especially in [Far+18, GR19].
@@
We now outline the “structural property” of a distribution that is at the core of all our proofs. Somewhat informally, it states that if $I(A:X|Y)$ and $I(A:Y|X)$ are sufficiently small, then a part of the mutual information between $X$ and $Y$ can be “materialized” in a sense analogous to the notion of Gács–Körner common information [GK73]. This means that there exists a random variable $W$ that captures a portion of the mutual information between $X$ and $Y$: on the one hand, this $W$ has negligible conditional entropy given either $X$ or $Y$; on the other hand, $W$ retains a significant amount of information: conditioned on $W$, the mutual information between $(X,Y)$ and $A$ becomes very small. The lemma below is our main technical tool.
Lemma D (name=,restate=lWkbounds).
@@
Suppose a tuple $(X,Y,A,B,\dots)$ is such that for some $\varepsilon>0$
$\max\left\{I(A:X|Y),I(A:Y|X)\right\}\leq\varepsilon\cdot I(X:Y)$
@@
Then there exists a sequence of $W_{i}$, $i\in\mathbb{N}_{0}$ such that
@@
Corollary E (name=,restate=corWbound).
Suppose a tuple $(X,Y,A,B,\dots)$ is such that for some $\varepsilon>0$
$\max\left\{I(A:X|Y),I(A:Y|X)\right\}\leq\varepsilon\cdot I(X:Y)$
Then there exists a $W$ such that
@@
Here an important clarification is in order. Lemma D is not literally valid for tuples of jointly distributed random variables $(X,Y,A,B,\dots)$. A version of the lemma formulated directly for random variables would necessarily be considerably more cumbersome:
“For every $\varepsilon^{\prime}>0$ for all sufficiently large $k$, for the scaling of the original distribution $(X,Y,A,B)$ obtained by taking $k$ independent, identically distributed copies of the variables, there exists a variable $W_{i}$ satisfying conditions (a), (b), and (c) within accuracy $\varepsilon^{\prime}\cdot k$.”
@@
Although this would technically suffice, it would obscure the structural features on which our analysis relies and complicate future integration of the lemma into automated proof frameworks. For this reason, Lemma D is stated not for distributions of random variables themselves, but for tropicalized random variables, which are abstract objects derived from ordinary random variables (see [MP20]). The definition and basic properties of tropicalized variables are discussed in Section 2.2. The tropicalization procedure provides a convenient “interface” that hides certain technical steps — such as scaling up the distribution and controlling sublinear slack terms — and thereby allows us to give a cleaner and more usable formulation of the technical Lemma D and Corollary E. By contrast, our main results (Theorems ABC, and F), which rely on Lemma D and Corollary E, can be understood in the more usual way: they apply directly to conventional tuples of random variables.
@@
We prove Lemma D and Corollary E in Section 3. The proofs may be viewed as compactly capturing the underlying construction originally due to Matúš. In particular, the proof of Corollary E requires an a priori unbounded number of applications of the Copy Lemma. This fact effectively encodes an infinite family of non-Shannon–type information inequalities, even though they never appear explicitly in the argument.

### v2

@@ 1 Introduction
In this paper we study inequalities for Shannon entropy, in particular variants of the Ingleton inequality in the entropic setting. At a high level, the goals of this paper can be summarized as two broad objectives:
  1. (I)
    gaining a better understanding of the geometry of the boundary of the cone of almost-entropic vectors;
  2. (II)
    developing tools for further progress in understanding conditional independence structures.
More specifically, we focus on conditional versions of the Ingleton inequality and on related structural properties of entropy. It is well known that the Ingleton inequality in its original form does not hold for the entropies of certain collections of jointly distributed random variables; see, e.g. [Mat95, Csi96, ZY98, Mat99, HRSV00]. However, under certain restrictions (or with suitable additional terms) the inequality may still hold. We study several manifestations of this phenomenon.
1.1 The cone of almost-entropic points
We begin with a brief review of the standard framework for information inequalities. For a tuple of random variables $(X_{i}\;{\bm{:}}\;i=0,\dots,n-1)$ its entropy profile is the function on the power set of the set $\left\{0,\dots,n-1\right\}$ indexing the tuple whose value on a set $J\in 2^{[n]}$ is the entropy of the joint distribution $(X_{i}\;{\bm{:}}\;i\in J)$. Since the entropy of the empty collection is always zero, we drop the corresponding value from consideration. Thus, every entropy profile is effectively a point in $\mathbb{R}^{2^{n}-1}$. In particular, for a four-tuple of random variables, its entropy profile is a 15-dimensional vector.
A point in $\mathbb{R}^{2^{n}-1}$ is called entropic if it is the entropy profile of the probability distribution of some tuple $(X_{i}\;{\bm{:}}\;i=0,\dots,n-1)$. The set of all entropy profiles for all $n$-tuples of random variable is usually denoted $\Gamma_{n}^{\star}$. The points in the closure of the set of entropy profiles are called almost-entropic points. This set is denoted by $\bar{\Gamma}_{n}^{\star}$. It has been shown in [ZY97] that this set is in fact a convex cone. The geometry of this cone is quite non-trivial. An exact description of this cone for $n\geq 4$ is still unknown. However, there are some known inner and outer bounds for this set. In what follows we briefly discuss the basic facts about $\bar{\Gamma}_{4}^{\star}$.
@@ 1.1 The cone of almost-entropic points
The set of all $\varphi$ satisfying these inequalities is traditionally denoted $\Gamma_{4}$ (the cone of submodular functions on a four-point set). This set is clearly a superset (i.e., an outer bound) of $\bar{\Gamma}_{4}^{\star}$. In other words, the axioms of polymatroids (non-negativity, monotonicity, and submodularity) are necessary conditions for a point to be entropic.
The cone $\bar{\Gamma}_{4}^{\star}$ is strictly smaller than $\Gamma_{4}$; that is, besides the inequalities valid for all polymatroids, there exist other linear inequalities that are valid for all almost-entropic points (for distributions of $n\geq 4$ random variables). The inequalities that are valid for all almost-entropic points but do not follow from the Shannon inequalities are called non-Shannon type inequalities. The first example of such an inequality was discovered by Z. Zhang and R. W. Yeung in [ZY98]. In the following years many other non-Shannon type inequalities were discovered; see, e.g., [MMRV02, DFZ06, XWS08, DFZ11, Csi16]. Moreover, F. Matúš proved in [Mat07b] that there exist infinitely many independent non-Shannon type inequalities (in other words, $\bar{\Gamma}_{4}^{\star}$ is not polyhedral, as well as all $\bar{\Gamma}_{n}^{\star}$ for $n>4$).
In the early years after the discovery of the first non-Shannon type inequalities, much effort was devoted to understanding how large the gap between $\bar{\Gamma}_{4}^{\star}$ and $\Gamma_{4}$ is. Accordingly, particular interest was placed on non-Shannon inequalities that exclude a significant region of the cone $\Gamma_{4}$. In recent years, attention has shifted to somewhat different questions. For example: what is the geometry of the boundary of the cone $\Gamma_{4}$? How accurately can it be approximated by piecewise algebraic surfaces? How smoothly does the curved part of the boundary approach a given flat face? The main results of this paper are primarily relevant to the second type of questions, namely those related to the geometry of the boundary of $\bar{\Gamma}_{4}^{\star}$ (see Objective (I) on p. (I)).
The main results of this paper are centered around the Ingleton inequality described in [Ing71]. To discuss this inequality, we introduce the following notation: for a set function $\varphi$ define
@@
Now the Ingleton inequality can be written in the following compact form:
@@
Permuting the set $\left\{0,1,2,3\right\}$ we obtain $\binom{4}{2}=6$ similar inequalities that differ in the order of the variables. We denote by $\Gamma^{\rm Ing}_{4}$ the cone of all non-negative, monotone, submodular functions $\varphi$ on $\left\{0,1,2,3\right\}$ that satisfy, in addition to the Shannon inequalities, the six versions of the Ingleton inequality on a four-point set (this cone under a different notation was discussed, e.g., in [CGK10, WC17]).
The Ingleton inequality is a sufficient condition for being an entropic point, i.e., every $\varphi$ in $\Gamma^{\rm Ing}_{4}$ is entropic, see [MS95] (independently, this was proved in [HRSV00]). In other words, $\Gamma^{\rm Ing}_{4}$ is a subset (inner bound) of the set of almost-entropic points. It is well known that the Ingleton inequality is not a necessary condition for being an entropic point, i.e., there exist entropic $\varphi$ outside $\Gamma^{\rm Ing}_{4}$, see [Mat95, Csi96, ZY98, Mat99, HRSV00, BN20, MC16].
@@
The sets $\Gamma_{4}$ and $\Gamma_{4}^{\mathrm{Ing}}$ are both polyhedral cones, and the set of almost-entropic points is sandwiched between the two: $\Gamma_{4}^{\mathrm{Ing}}\subsetneq\bar{\Gamma}_{4}^{\star}\subsetneq\Gamma_{4}.$ The closure of $\Gamma_{4}\setminus\Gamma_{4}^{\mathrm{Ing}}$ consists of six closed convex cones with disjoint interiors, [MP20]. They are symmetric under permutations of the set $\left\{0,1,2,3\right\}$. From now on, we consider one of these cones in which the strict inequality (1) is violated and, following [MP20], refer to it as the non-Ingleton cone. Let us briefly recall how this object can be viewed in geometric terms (see [MP20] or [CGK10, GCG11] for more details). The non-Ingleton cone has a 14-dimensional simplex as its base. One of the 13-dimensional faces of this simplex is defined by the Ingleton equality and is a common face with the base of $\Gamma_{4}^{\mathrm{Ing}}$. The remaining vertex of the simplex is the so-called special point. It is neither an entropic nor an almost entropic point, as was shown in [ZY98].
@@
The study of the cone of almost-entropic points essentially reduces to the question of what neighborhood around the special point is free of almost entropic points, or, equivalently, what neighborhood of the face shared between the Ingleton and non-Ingleton regions is filled with almost entropic points. [^1] If we denote by $\alpha_{1},\dots,\alpha_{15}$ the convex coordinates in the non-Ingleton simplex (the base simplex of non-Ingleton cone), such that $\alpha_{15}=1$ is the defining equation of the special point, while $\alpha_{15}=0$ is the equation of the face common with the Ingleton-region, then the almost -entropic region is described by the inequality $\alpha_{15}\leq\Phi(\alpha_{1},\dots,\alpha_{14})$ where $\Phi$ is some concave function on the $13$-dimensional simplex. Ultimately, the complete characterization of $\bar{\Gamma}^{*}_{4}$ reduces to the description of the function $\Phi$. The main results of this paper contribute to a better understanding of the behavior of $\Phi$ near the boundary of its domain. Thus, the Ingleton inequality plays a crucial role in the geometry of $\bar{\Gamma}^{*}_{4}$. On the one hand, this inequality together with all its permutations provides an inner bound for this cone. On the other hand, many non-Shannon type inequalities can be written in the form of the Ingleton inequality with a relatively small additional “correction term”, which give an outer bound for this cone, see the discussion in [WC17, section 2]. Systematic investigations of the geometry of the cones $\Gamma_{4}$, $\bar{\Gamma}_{4}^{\star}$, and $\Gamma_{4}^{\mathrm{Ing}}$ can be found in [Csi25, MC16, MP20]. In general, the complete description of $\bar{\Gamma}_{4}^{\star}$ remains far from being understood, despite considerable efforts devoted to this problem.
1.2 Constraint versions of the Ingleton inequality.
@@
In what follows we consider the Ingleton inequality in the entropic setting and discuss several versions of this inequality obtained by imposing constraints and/or adding a negligible correction term, that are valid for entropic and almost-entropic points. To avoid ambiguity and improve the readability of the notation we denote the four-tuple of jointly distributed random variables $(X,Y,A,B)$, and rewrite (1) with the usual notation for the conditional mutual information as
@@
When discussing constraint inequalities, we assume that some of the quantities
I(X:A|Y),I(Y:A|X),I(X:B|Y),I(Y:B|X),I(X:Y|A),I(X:Y|B)I(X:A|Y),I(Y:A|X),I(X:B|Y),I(Y:B|X),I(X:Y|A),I(X:Y|B) (2)
are small compared to the other information quantities. It is known that (ING) holds for joint distributions $(X,Y,A,B)$ satisfying certain independence constraints, that is, whenever some of the quantities in (2) vanish, see [KR13, Stu21]. We will focus on the following two implications:
@@
To the best of our knowledge, the conditional inequalities (3) and (4) were first explicitly formulated in this form in [KR13]. The proof given in [KR13] relied on a deep and nontrivial result of Matúš, namely the infinite series of conditional inequalities established in [Mat07b]. Studený proposed in [Stu21] a different proof of (3) and (4), which is more intuitive and technically simpler. His argument is based on the following simple lemma.
Lemma 1 (see lemma 8 in [Stu21]).
Every discrete distribution of random variables $(X,Y,A)$ satisfying
I(X:A|Y)=0,I(Y:A|X)=0I(X:A|Y)=0,\ I(Y:A|X)=0
can be extended by a discrete random variable $W$ such that in the joint distribution $(X,Y,A,W)$ the variable $W$ is functionally dependent on both $X$ and $Y$, and moreover
$I(X,Y:A|W)=0.$
The proof of (3) and (4) then proceeds by applying this lemma in combination with standard Shannon inequalities. Conceptually, this argument is reminiscent of the classical proof of the Ingleton inequality for linearly representable matroids.
Remark 2.
Lemma 1 and its use in proving information inequalities are analogous to the Ahlswede–Körner lemma and its role in the derivation of non-Shannon-type information inequalities in [MMRV02]. Recall that the Ahlswede–Körner lemma was used in [FKMP18] in computer-assisted proofs of lower bounds for problems of secret sharing. We believe that Lemma 1 (and its “smooth” versions discussed below) can be also used in computer-assisted proofs of conditional information inequalities.
However, the simplicity of Studený’s approach comes at a price. Lemma 1 was proved in [Stu21] for entropic points, but not for almost-entropic ones. Consequently, the resulting proof of (3) and (4) applies only to entropic points (whereas the more involved argument in [KR13] shows that these inequalities also hold for almost-entropic points).
In this paper we introduce a technical tool that combines the transparency of Studený’s approach with the power of Matúš’s method. We prove a robust version of Lemma 1 that applies to almost-entropic points, see Corollary F and Corollary G. Moreover, we establish a trade-off between the error in the assumptions
$I(X:A|Y)\approx 0,\qquad I(Y:A|X)\approx 0$ (5)
and the possible deviation in the conclusions
$H(W|X)\approx 0,\qquad H(W|Y)\approx 0,\qquad I(XY:A|W)\approx 0.$ (6)
That is, we assume that the approximate equations in (5) hold with precision $\varepsilon$ and analyze the resulting asymptotics in (6).
Throughout the paper we formulate results of this type using asymptotic notation, without specifying the multiplicative constants hidden in the $O(\cdot)$ terms. We believe that refining the asymptotic behavior, rather than optimizing specific constants, will help reveal fundamental properties of the cone of almost-entropic points, see also Remark 3. For a discussion of non-asymptotic and unconditional versions of Theorems A, B, and C, see Remark 4.
Let us mention that (3) and (4) helped to classify the conditional independence structures over $4$ random variables, [Stu21, Boe24] (cf. Objective (II) on p. (II)). The classification of similar structures for $n>4$ remains an open problem, partly due to the rapidly increasing technical difficulties as the dimension grows. We hope that the conceptual tools proposed in this paper, such as the robust version of Lemma 1, may contribute to progress in this direction.
1.3 Main technical results
We focus on the possible violation of the Ingleton inequality in the regime where some of the mutual information terms in (2) are relatively small but still positive. We consider three cases:
@@ 1.3 Main technical results
$\max\left\{I(X:A|Y),I(Y:A|X)\right\}\leq\varepsilon\cdot I(X:Y).$
@@
$\max\left\{I(Y:A|X),I(X:Y|A)\right\}\leq\varepsilon\cdot I(X:A).$
@@
These three theorems capture certain geometric features of the boundary of the cone $\bar{\Gamma}^{\star}_{4}$, in particular on how smoothly the curved part of the boundary approaches the flat face defined by (2).
@@
$\quad{}+\frac{1}{k}I(X:Y|A)+\frac{k+1}{2}\big(I(X:A|Y)+I(Y:A|X)\big)\quad\text{for}\ k\in\mathbb{N}$
@@
$\quad+\frac{1}{k}I(Y:A|X)+\frac{k+1}{2}\big(I(X:Y|A)+I(X:A|Y)\big)\quad\text{for }k\in\mathbb{N}$
proven by Matúš in [Mat07b] (it suffices to choose a suitable value of $k$). In fact, our proof of the lemma above runs parallel to the argument of Matúš — we apply the Copy Lemma and the Ahlswede–Körner lemma, while Matúš used in [Mat07b] a similar (and formally equivalent) method of adhesive extensions.
@@
In this paper we present more direct proofs of Theorems AB, and C without deducing preliminary infinite families of non-Shannon–type inequalities for entropy.
Another result proved in this paper reveals more subtle properties of probability distributions. It can be presented as the following alternative. If in a six-tuple of random variables $(X,Y,A,B,C,D)$ the four quantities
I(X:A|Y),I(Y:A|X),I(X:B|Y),I(Y:B|X)I(X:A|Y),\ I(Y:A|X),\ I(X:B|Y),\ I(Y:B|X)
are negligibly small, then one of the two properties must be true: a stronger version of the Ingleton inequality for $(X,Y,A,B)$,
$I(X:Y)\textrm{ is much less than }I(X:Y|A)+I(X:Y|B)+I(A:B),$
or a slightly weaker version of the Ingleton inequality for $(X,Y,C,D)$,
$I(X:Y)\ \textrm{ is not much bigger than }I(X:Y|C)+I(X:Y|D)+I(C:D).$
Here is the exact statement.
Theorem D (name=,restate=thsubtler).
If the joint distribution $(X,Y,A,B,C,D)$ satisfies
$\max\{I(X:A|Y),I(Y:A|X),I(X:B|Y),I(Y:B|X)\}\leq\varepsilon\cdot I(X:Y)$
Then
  1. (a)
    The following alternative holds: for every threshold $0<t<1$ we have a stronger version of Ingleton’s inequality for $(A,B,X,Y)$
    or a weaker version of Ingleton’s inequality for $(X,Y,C,D)$
    $I(X:Y)\leq I(X:Y|C)+I(X:Y|D)+I(C:D)+\big(O(\sqrt{\varepsilon})+t\big)I(X:Y)$
  2. (b)
    Besides, we have the following inequality for six random variables
    $\;\;+I(X:Y|C)+I(X:Y|D)+I(C:D)$
Note that the assumptions in the theorem above do not involve variables $C,D$ in any capacity, so that the conclusion of the theorem remains true for arbitrary $C$, $D$.
While Theorems AB, and C admit proofs based on explicitly derived infinite families of non-Shannon-type inequalities, for Theorem D (proved in Section 5) we are are not aware of any alternative proof that differs substantially from our approach.
In the next section we give a brief overview of our technique.
1.4 Methods and ideas
@@
This article presents more direct proofs of Theorems AB (Section 4), and Theorem C (Section 6) without deducing preliminary infinite families of non-Shannon–type inequalities for entropy. This approach offers, in our view, several noteworthy advantages. To begin with, it is conceptually more intuitive, as it is grounded in certain structural properties of probability distributions and in the intuition of materializable common information, which we elaborate on below. This perspective enables us to reveal and prove properties that are stronger than a mere conditional form of the Ingleton inequality, as in Theorem D. In addition, the structural property we identify appears particularly amenable to computer-assisted proof systems. We believe that this may enable computer-assisted discoveries of new conditional information inequalities, in much the same spirit as non-Shannon–type inequalities and their applications have been established in [DFZ11] and especially in [FKMP18, GR19].
@@
We now outline the structural property of a distribution that is at the core of all our proofs. Somewhat informally, it states that if $I(X:A|Y)$ and $I(Y:A|X)$ are sufficiently small, then a part of the mutual information between $X$ and $Y$ can be “materialized” in a sense analogous to the notion of Gács–Körner common information [GK73]. This means that there exists a random variable $W$ that captures a portion of the mutual information between $X$ and $Y$. On the one hand, this $W$ has negligible conditional entropy given either $X$ or $Y$; on the other hand, $W$ retains a significant amount of information—conditioned on $W$, the mutual information between $(X,Y)$ and $A$ becomes very small. This idea is formalized in Lemma E below, which is our main technical tool.
Before Lemma E is formulated, we need to introduce two notions: symmetry among random variables and tropical probability spaces. We begin with the symmetry. Suppose that a tuple $(X,Y,A,B,Z)$ is such that the variables $A$ and $B$ take value in the same alphabet and such that the joint distribution $p$ of the tuple satisfies
$p(x,y,a,b,z)=p(x,y,b,a,z)$
for all samples $x,y,z$ in the alphabets of $X,Y,Z$, respectively, and all $a,b$ in the (shared) alphabet of $A$ and $B$. In that case we say that the tuple is symmetric with respect to transposing $A$ and $B$ or variables $A$ and $B$ are symmetric within the tuple.
Tropical probability spaces were introduced in [MP18]. Tropicalization associates to a random variable $X$ (or a tuple of random variables) a sequence $X^{tr}:=(X^{(n)})_{n\in\mathbb{N}_{0}}$, whose elements behave, in an appropriate sense, like the joint distribution composed of $n$ independent copies of $X$. The entropy of such a tropical random variable is defined as
$H(X^{tr}):=\lim_{n\to\infty}\tfrac{1}{n}H\big(X^{(n)}\big)$
with the obvious generalization for tuples and entropic expressions on tuples. This is similar to the tensorization trick, which is widely used in analysis, geometry, probability, and related areas. In Section 2.2 we discuss the technique of tropicalization in more detail.
One of the advantages of tropicalization is that any relation between entropic quantities that holds for classical random variables up to a sublinear error holds exactly in the tropical setting. From now on we omit the superscript $tr$ and simply refer to such objects as tropical. Properties of tuples are understood element-wise along the tropical sequence. For example, a tropical pair $(X,Y)$ is called symmetric if the pair $\big(X^{(n)},Y^{(n)}\big)$ is symmetric for all sufficiently large $n$.
Lemma E (name=,restate=lWkbounds).
@@
Suppose a tropical tuple $(X,Y,A,B,\dots)$ is such that for some $\varepsilon>0$
$\max\left\{I(X:A|Y),I(Y:A|X)\right\}\leq\varepsilon\cdot I(X:Y)$
@@
Then there exists a sequence of tropical random variables $W_{i}$, $i\in\mathbb{N}_{0}$ such that
@@
Corollary F (name=,restate=corWbound).
Suppose a tropical tuple $(X,Y,A,B,\dots)$ is such that for some $\varepsilon>0$
$\max\left\{I(X:A|Y),I(Y:A|X)\right\}\leq\varepsilon\cdot I(X:Y)$
Then there exists a tropical $W$ such that
@@
This implies immediately another corollary:
Corollary G.
Lemma 1 holds for almost-entropic points.
@@
The tropicalization procedure provides a convenient “interface” that hides certain technical steps — such as scaling up the distribution and controlling sublinear slack terms — and thereby allows us to give a cleaner and more usable formulation of the technical Lemma E and Corollary F. By contrast, our main results (Theorems ABC, and D), which rely on Lemma E and Corollary F, can be understood in the more usual way: they apply directly to conventional tuples of random variables.
@@
We prove Lemma E and Corollary F in Section 3. The proofs may be viewed as compactly capturing the underlying construction originally due to Matúš. In particular, the proof of Corollary F requires an a priori unbounded number of applications of the Copy Lemma. This fact effectively encodes an infinite family of non-Shannon–type information inequalities, even though they never appear explicitly in the argument.